OpenStudy (anonymous):

What is the value of x in the solution to the system of equations below? 15x – 12y = 13 30x + 9y = 4 A. -17/3 B. 1/3 C. -2/3 D. 1/6

OpenStudy (anonymous):

First you have to choose a path to solve the equation system There're some basic ways to do that, you can isolate one variable in an equation, for example in the first eq i'll isolate "y" 15x-12y=13 /* +12y-13 */ 15x-12y+12y-13=13+12y-13 15x-13=12y /* (*1/2) */ y= (15x+13)/12 And then replace this value of y in the other eq. Give it a try